# How Grounded is Wikipedia? A Study on Structured Evidential Support

### v1

@@ Abstract
Wikipedia is a critical resource for modern NLP, serving as a rich repository of up-to-date and citation-backed information on a wide variety of subjects. The reliability of Wikipedia—its groundedness in its cited sources—is vital to this purpose. This work provides a quantitative analysis of the extent to which Wikipedia is so grounded and of how readily grounding evidence may be retrieved. To this end, we introduce PeopleProfiles  [^1] https://github.com/wgantt/people-profiles—a large-scale, multi-level dataset of claim support annotations on Wikipedia articles of notable people. We show that $\sim 20\%$ of claims in Wikipedia lead sections are unsupported by the article body; $\sim 27\%$ of annotated claims in the article body are unsupported by their (publicly accessible) cited sources; and $>80\%$ of lead claims cannot be traced to these sources via annotated body evidence. Further, we show that recovery of complex grounding evidence for claims that are supported remains a challenge for standard retrieval methods.
@@ 1 Introduction
Long an essential ingredient for LLM pretraining, Wikipedia is now widely used during inference as a repository of high-quality, citation-backed information for RAG applications (Lewis et al., 2020; Chen et al., 2020; Fan et al., 2024, i.a.). In parallel, Wikipedia has played a major role in advancing automated fact or claim verification (Dmonte et al., 2024), enabling the creation of many notable benchmarks for these tasks, such as FEVER (Thorne et al., 2018a, b), WikiFactCheck-English (Sathe et al., 2020), VitaminC (Schuster et al., 2021), and WICE (Kamoi et al., 2023). However, whereas these works treat Wikipedia articles as sets of claims or passages to sample from for dataset curation, this work studies Wikipedia articles as whole, structured documents—relied upon as trustworthy sources for information-seeking tasks. [^2] Of these, WICE is most similar to our work. Appendix C has a detailed discussion of differences.
First, we ask to what extent claims in Wikipedia are grounded. Acknowledging Wikipedia’s distinction between an article’s lead section (i.e. the intro) and its body, we are the first to jointly explore both how claims in the lead are grounded in the body (article-internal support) and how claims in the body are in turn grounded in publicly accessible cited sources (article-external support). Second, we ask how effectively standard retrieval methods can recover evidence about these claims—either from the body (for lead claims) or from cited source documents (for body claims). In answering these questions, we make the following contributions:
@@
  • We show that: (1) a surprising proportion of lead claims ($\sim 20\%$) are unsupported by the body of the same article; (2) an even higher proportion of body claims ($\sim 27\%$) are unsupported by scrapable sources.; and (3) $>80\%$ lead claims cannot be directly grounded in these sources via supporting body sentences.
  • We show that even in Wikipedia, evidence for these claims is often complex, involving multiple premises, and that retrieval of such evidence remains challenging.
Figure 1: An example of the multi-level structure of PeopleProfiles annotations. Claims in the lead of a Wikipedia article (top left) are supported by sentences in the body (bottom left), whose claims in turn are supported by evidence in cited sources (right). Prior work on Wikipedia claim verification has not considered this structure.
@@ 2 Data Collection
Methodology
We obtain evidence for Wikipedia claims and scalar [-1,1] judgments of the degree of support/refutation for those claims given that evidence. [^3] While refutation is unlikely in Wikipedia, we wanted to capture the rare cases where it occurs. We divide annotation into two phases—one for claims appearing in the article’s lead and a second for claims appearing in its body. This is motivated by the different guidelines Wikipedia establishes for these two parts of an article: while citations are required for contentful claims in the body (e.g. quotations, statistics), [^4] https://en.wikipedia.org/wiki/Wikipedia:When_to_cite “it is common for citations to appear in the body and not the lead,” since “significant information should not appear in the lead if it is not covered in the remainder of the article.” [^5] https://en.wikipedia.org/wiki/Wikipedia:Manual_of_Style/Lead_section Thus, for lead claims, we seek evidence in the body, and for body claims, we seek evidence in cited sources. Following prior work (Kamoi et al., 2023), we define the evidence for a claim as a set of (possibly non-contiguous) sentences. We annotate up to 3 sentences that together provide the strongest evidence for or against each target claim.
Claims
We adopt the view advocated in work on claim decomposition that the appropriate units for assessment of evidential support are subclaims, i.e., sub-sentence-level statements expressing an atomic proposition (Kamoi et al., 2023; Min et al., 2023; Wanner et al., 2024a, b; Gunjal and Durrett, 2024, i.a.). [^6] When we refer to claims in this work, we mean subclaims. We use the “DnD” method of Wanner et al. (2024b) to jointly decompose each Wikipedia sentence into two sets of subclaims: a contextualized set decomposed from the sentence alone and a decontextualized set that inserts into each subclaim relevant extra-sentential context (e.g. to resolve pronouns). Following Wanner et al., we use GPT-4o-mini (OpenAI, 2024) to perform the decomposition. [^7] See Appendix A for prompts. Annotators can see both versions of a subclaim when assessing its support.
@@
Data Source
@@
We select for annotation Wikipedia articles of notable people (entities) studied in prior work on claim verification, including all those from Min et al. (2023) and Jiang et al. (2024), yielding 1,485 entities that represent a range of nationalities and degrees of fame. We obtain the (structured) English articles for each entity from the MegaWika 2.0 dataset (Barham et al., 2023, 2025), including their in-text citations and the citations’ scraped source texts. We annotate claim support for subclaims decomposed from all sentences in articles’ leads and all body sentences that bear citations to publicly accessible sources, as we cannot verify paywalled or print sources at scale—nor can Wikipedia users or RAG-enabled search engines. We use the DeBERTa-based (He et al., 2020) text quality classifier from NVIDIA’s NeMo Curator to filter low-quality sources. [^8] https://github.com/NVIDIA/NeMo-Curator We divide entities into train (65%) and dev and test (35%) splits via stratified sampling on the number of lead subclaims.
Pilot Annotation
To ensure high-quality automatic annotation on the full entity set, we conduct a pilot human annotation on a set of 160 body claims obtained from 10 entities, divided into 3 batches. Each batch was annotated with two-way redundancy by three authors, using an interface and instructions we designed for the task (see Appendix A). We assess inter-annotator agreement on support judgments using Krippendorff’s $\alpha$ (Krippendorff, 2018) and on the selected sets of evidence sentences using average pairwise $\text{F}_{1}$ (with exact sentence match), obtaining $\alpha=54.3$ and $\text{F}_{1}=53.8$. We then use these results to guide prompt engineering for the bulk annotation, assessing GPT-4o-mini on the same examples. Model annotations from our final (best) prompt yield $\alpha=64.6$ and $\text{F}_{1}=54.1$ when included with the original human ones—indicating that GPT-4o-mini achieves inter-human agreement levels.
Bulk Annotation
@@
Using GPT-4o-mini with the same prompt, we collect support and evidence annotations on all 1,485 entities. Table 1 shows statistics of the resulting PeopleProfiles dataset.
@@
3 Analysis & Experiments
3.1 Claim Support
A significant fraction of lead and body claims are unsupported.
Figure 2 plots lead and body claim support distributions for the PeopleProfiles dev split. We observe strong bimodality in both distributions, with high density around both full support (1.0) and no support (0.0). Indeed, 19.3% of lead claims are judged unsupported ($\leq 0$) by the body text and 26.5% of body claims by their cited source text(s). Inspection reveals that, contrary to guidelines, many leads make assertions attested nowhere else in the article—notably, about birth and death date and location—while other unsupported claims present inherently difficult attribution problems (e.g. nickname origins). Similarly, many body claims assert propositions unattested in publicly accessible sources: numerous articles extensively cite copyrighted books or paywalled articles which, while clearly legitimate, places hard limits on the amount of content that can be readily verified by (human or machine) readers.
Support does not robustly propagate from sources up to lead claims.
@@ 3.1 Claim Support
We directly annotate lead claim support given body evidence, but we can further consider how strong the support is for that evidence based on cited sources. We consider two methods of computing a support score for a body evidence sentence given the support scores of its decomposed claims, either taking the scores’ mean or their product (clipping scores $<0$ to 00). We can then compute an overall score for an evidence set via the same aggregations applied to the set, yielding 4 possible overall scores. Broadly we find that (1) most lead claims (82%) do not ground out in source evidence because their body evidence sentences lack citations; (2) of those that do, average overall evidence scores are very modest (e.g. 0.41 when using mean-mean aggregation). [^9] Appendix B plots these overall score distributions.
Task Model N@5 R@5 R@10
LEAD ColBERTv2 52.59 57.90 68.18
Stella-1.5B-v5 30.03 38.03 51.35
BM25 49.92 56.02 66.01
+Rerank Rank1 60.55 63.25 66.01
BODY ColBERTv2 70.02 76.37 87.16
Stella-1.5B-v5 49.37 60.89 78.15
BM25 61.70 68.21 80.24
+Rerank Rank1 73.02 76.18 80.24
Task Model N@5 N@100 R@100
ENTITY ColBERTv2 24.53 39.65 62.52
Stella-1.5B-v5 13.56 28.40 57.96
BM25 14.59 30.08 56.67
+Rerank Rank1 20.54 34.28 56.67
Table 2: Evidence retrieval results for LEAD, BODY, and ENTITY. Best first-stage results are bolded. “+Rerank” is reranked BM25 results ($k=10$ for LEAD and BODY; $k=100$ for ENTITY). N=NDCG; R=Recall.
3.2 Evidence Retrieval
We consider three evidence retrieval tasks:

### v2

@@ Abstract
Wikipedia is a critical resource for modern NLP, serving as a rich repository of up-to-date and citation-backed information on a wide variety of subjects. The reliability of Wikipedia—its groundedness in its cited sources—is vital to this purpose. This work analyzes both how grounded Wikipedia is and how readily fine-grained grounding evidence can be retrieved. To this end, we introduce PeopleProfiles  [^1] https://github.com/wgantt/people-profiles—a large-scale, multi-level dataset of claim support annotations on biographical Wikipedia articles. We show that: (1) ${\sim}22\%$ of claims in Wikipedia lead sections are unsupported by the article body; (2) ${\sim}30\%$ of claims in the article body are unsupported by their publicly accessible sources; and (3) real-world Wikipedia citation practices often differ from documented standards. Finally, we show that complex evidence retrieval remains a challenge—even for recent reasoning rerankers.
@@ 1 Introduction
Long an essential ingredient for LLM pretraining, Wikipedia is now widely used during inference as a repository of high-quality, citation-backed information for RAG applications (Lewis et al., 2020; Chen et al., 2020; Fan et al., 2024, i.a.). In parallel, Wikipedia has played a major role in advancing automated fact or claim verification (Dmonte et al., 2024), enabling the creation of many notable benchmarks for these tasks, such as FEVER (Thorne et al., 2018a, b), WikiFactCheck-English (Sathe et al., 2020), VitaminC (Schuster et al., 2021), and WiCE (Kamoi et al., 2023). But whereas these works treat Wikipedia articles as sets of claims or passages to sample from for dataset curation, this work studies Wikipedia articles as whole, structured documents—relied upon as trustworthy sources for information-seeking tasks.
First, we ask how grounded claims in Wikipedia are. Acknowledging Wikipedia’s distinction between an article’s lead section (i.e. the intro) and its body (the remaining sections), we are the first to jointly explore both how claims in the lead are grounded in the body (article-internal support; Figure 1, top arrow) and how claims in the body are in turn grounded in publicly accessible cited sources (article-external support; Figure 1, bottom arrow). Second, we ask how well both standard first-stage retrievers and new reasoning-based reranking models can recover evidence about these claims—either from the body (for lead claims) or from cited source documents (for body claims). In answering these questions, we make the following contributions:
Figure 1: PeopleProfiles features fine-grained, multi-level evidential relations and scalar support labels (not shown) on Wikipedia articles—from cited sources to claims in the article body (bottom arrow), and from the body to claims in the article lead (top arrow).
@@
  • We show that: (1) a surprising proportion of lead claims (${\sim}22\%$) are unsupported by the body of the same article; (2) an even higher proportion of body claims (${\sim}30\%$) are unsupported by scrapable sources; and, more generally, (3) actual Wikipedia citation practices often differ from documented standards.
  • We show that evidence for these claims is often complex, involving multiple premises, and that retrieval of such evidence remains challenging even for new reasoning rerankers.
Figure 2: A more detailed view of the multi-level structure of PeopleProfiles annotations. Claims in the lead of a Wikipedia article (top left) are supported by sentences in the body (bottom left), whose claims in turn are supported by evidence in cited sources (right). Prior work on Wikipedia claim verification has not considered this structure.
2 Background
Wikipedia-Based Claim Verification
Fact or claim verification is an active area of research within NLP with many supporting datasets, including several that draw heavily on Wikipedia. [^2] We refer the reader to Dmonte et al. (2024) for a general survey of claim verification and focus only on Wikipedia here. Of these, FEVER (F; Thorne et al., 2018a, b) is the most widely studied, featuring 185K claims written by human annotators, given randomly chosen Wikipedia sentences as prompts. For each claim, a different set of annotators then curated evidence passages—sets of sentences from one or more Wikipedia pages—as well as claim support labels (Supported, Refuted, NotEnoughInfo).
WikiFactCheck-English (WFC; Sathe et al., 2020) consists of 125K claims also derived from Wikipedia, but uses binary (Supported, Refuted) labels. Supported claims are Wikipedia sentences with at least one citation and Refuted claims are obtained via manual edits to the original (true) claims to make them false. Evidence consists of full documents cited by the claim sentence.
VitaminC (VC; Schuster et al., 2021) uses factual revisions made by Wikipedia editors on a large set of articles as well as synthetic revisions to evidence sentences from F to build a set of contrastive evidence sentence pairs, based on which annotators then write Supported or Refuted claims—yielding 326K claim-evidence pairs.
Finally, the work most similar to ours is WiCE (W; Kamoi et al., 2023), which annotates ternary support labels on subclaims automatically decomposed from Wikipedia sentences drawn from the SIDE dataset (Petroni et al., 2023). [^3] Table 8 summarizes key differences between our work and W. Differences with F/WFC/VC are discussed in §2. Each subclaim is paired with annotated sets of evidence sentences from associated cited web articles. In total, WiCE contains roughly 5.4K subclaim-evidence pairs.
Claim Decomposition
A wealth of recent work on claim decomposition has argued that the appropriate units for assessment of evidential support are subclaims, i.e., sub-sentence-level propositions (Kamoi et al., 2023; Min et al., 2023; Wanner et al., 2024a, b; Gunjal and Durrett, 2024, i.a.). Broadly, this argument contends that because subclaims assert single atomic propositions, they are easier and less ambiguous to evaluate than sentences, which may assert (or presuppose) multiple propositions. In practice, the optimal conception of atomicity is contested, and a variety of decomposition techniques have been proposed (Wanner et al., 2024b; Gunjal and Durrett, 2024). While we do not contribute to this debate, we accept the consensus that undergirds it—that subclaims are to be preferred to sentences—leveraging an existing claim decomposition method to construct PeopleProfiles3).
Our Work
Our PeopleProfiles dataset differs from the resources discussed above in several key ways. First, our claims are ecologically valid, as they are sourced from Wikipedia articles, and not synthetically constructed or perturbed by crowdworkers (contra VC, WFC). Second, we annotate support on atomic subclaims, not full sentences, making it unambiguous which proposition’s support is being assessed (contra F, VC, WFC). Third, we obtain scalar (not categorical) support judgments to capture variation in the degree of partial support that subclaims may have. Finally, our annotations capture multiple levels of evidence—enabling us to trace support from key claims in Wikipedia lead sections down into the body, and from there into cited sources. These last two features are unique to PeopleProfiles.
@@ 3 Data Collection
3.1 Methodology
We obtain evidence for Wikipedia claims and scalar [-1,1] judgments of the degree of support/refutation for those claims given that evidence. [^4] While refutation is uncommon in Wikipedia, we wanted to be able to capture the cases where it occurs. We divide annotation into two phases—one for claims appearing in the article’s lead and a second for claims appearing in its body. This is motivated by the different guidelines Wikipedia establishes for these two parts of an article: while citations are required for contentful claims in the body (e.g. quotations, statistics), [^5] See Wikipedia:When_to_cite “it is common for citations to appear in the body and not the lead,” since “significant information should not appear in the lead if it is not covered in the remainder of the article.” [^6] See Wikipedia:Manual_of_Style/Lead_section Thus, for lead claims, we seek evidence in the body, and for body claims, we seek evidence in cited sources. Following prior work (Kamoi et al., 2023), we define the evidence for a claim as a set of (possibly non-contiguous) sentences. We annotate up to three sentences that together provide the strongest evidence for or against each target claim.
3.2 Claims
Per §2, we adopt the view advocated in work on claim decomposition that the appropriate units for assessment of evidential support are subclaims, i.e., sub-sentence-level statements expressing an atomic proposition (Kamoi et al., 2023; Min et al., 2023; Wanner et al., 2024a, b; Gunjal and Durrett, 2024, i.a.). [^7] When we refer to claims in this work, we mean subclaims in this sense. (The claim decomposition literature often uses claim synonymously with sentence.) We use the “DnD” method of Wanner et al. (2024b) to jointly decompose each Wikipedia sentence into two sets of subclaims: a contextualized set decomposed from the sentence alone and a decontextualized set that inserts into each subclaim relevant extra-sentential context that may help resolve ambiguities in the contextualized version (e.g. unresolved pronouns). Annotators annotate all subclaims decomposed from a target sentence and can toggle between the contextualized and decontextualized versions of a subclaim when identifying evidence and assessing support. Following Wanner et al. (2024b), we perform the decomposition using GPT-4o-mini (OpenAI, 2024). [^8] See Appendix A for prompts and details of DnD.
@@ 3.2 Claims
3.3 Data Source
Much prior work on claim decomposition and verification has focused on biographies of notable people—entities—as a natural test domain in which the ground truth (e.g. birth and death dates, educational background) is generally uncontroversial and where sources of evidence are numerous (Min et al., 2023; Jiang et al., 2024; Gunjal and Durrett, 2024). Given this, biographical Wikipedia pages also likely provide a rough upper bound on the degree of evidential support for claims in Wikipedia articles from other domains, which may be subject to wider debate or disagreement (e.g. theoretical frameworks, religions, or political movements). This upper bound would thus strengthen findings of a lack of evidential support in these articles (§4).
@@ 3.3 Data Source
Accordingly, we follow this line of prior work and annotate the Wikipedia pages for all 1,485 entities studied by Min et al. (2023) and Jiang et al. (2024), representing a wide range of nationalities and degrees of fame. We obtain the English articles for each entity from the MegaWika 2 dataset (Barham et al., 2025), which includes (inter alia) section boundaries, in-text citations, and citations’ scraped source texts for each article. We annotate sets of evidence sentences and scalar support labels for subclaims decomposed from (1) all sentences in articles’ leads and (2) all body sentences that bear citations to publicly accessible sources (whose contents are provided in MegaWika 2). We focus on public sources because neither we nor Wikipedia users—nor even RAG-enabled search engines—have the resources or licenses required to verify paywalled or print sources at scale.
Barham et al. (2025) acknowledge rare errors in their source scraping process for MegaWika 2, resulting in 404 responses or other incomplete rendering of source content. To mitigate this, we further filter out low-quality sources (e.g. error pages) using the DeBERTa-based (He et al., 2020) text quality classifier from NVIDIA’s NeMo Curator. [^9] https://github.com/NVIDIA/NeMo-Curator
Finally, we construct train (965 entities), dev (256), and test (264) splits via stratified sampling of entities based on their number of lead subclaims.
3.4 Pilot Annotation
Given the large scale of our (main) bulk annotation (${\sim}150$K claims; §3.5), we adopt an automatic annotation process. To verify the quality of this process, we first conduct a pilot human annotation on a collection of 160 body claims obtained from 10 entities, divided into 3 batches. Three authors participated in the pilot, with a unique pair of annotators doubly annotating each batch using an interface and instructions that we developed. [^10] See Appendix A for interface details, annotation instructions, and annotator demographics.
We then use these annotations to guide manual prompt engineering for the bulk annotation, assessing GPT-4o-mini on the same examples and aiming to optimize agreement with the human annotations along two dimensions: (1) average pairwise $\text{F}_{1}$ with annotated evidence sets (using exact sentence match as the underlying similarity); and (2) Krippendorff’s $\alpha$ (with interval difference function; Krippendorff, 2018) on the scalar support labels. Model annotations from our final (best) prompt on the pilot claims yield average pairwise $\alpha{=}66.3$ and $\text{F}_{1}{=}62.7$ w.r.t. human annotations. This label agreement is only slightly below human expert agreement reported for related works, including F ($\alpha{=}68.4$) and VC ($\alpha{=}70.7$), despite our use of scalar labels and allowance of deviation in the underlying evidence sets. [^11] These other works report agreement as Fleiss’s $\kappa$, which is a special case of $\alpha$. Agreement on evidence selection is also on par with human expert agreement on these same examples ($\text{F}_{1}{=}61.4$).
3.5 Bulk Annotation
@@
Using GPT-4o-mini with the best prompt identified in the pilot (same for lead and body claims), we collect support and evidence annotations on all 1,485 entities. Table 1 shows statistics of the resulting PeopleProfiles dataset. Table 8 further illustrates that our automatic pipeline enables annotation of roughly 20x more body claims than W (Kamoi et al., 2023), the most similar prior work.
@@
4 Claim Support
Claim Type Category Example Claim
LEAD Birth/death date/location Josh Mansour was born in 1990.
Nicknames Michael Barakan is known as Shane Fontayne.
Nationality David Thomas Broughton is English.
Career Status Richie Dorman is retired.
BODY Background/Summary Claims Yorktown proved to be the last campaign of the Revolutionary War.
Hearsay/Disputed Information Some accounts report that Washington opposed flogging.
Bleached Claims There was a cabinet.
Table 2: Some common categories of lead and body claims among those with an annotated support score $\leq 0$.
In this section, we present some analysis of Wikipedia claim support enabled by PeopleProfiles, including the degree of support for both body and lead claims, the kinds of claims that tend to be unsupported, and propagation of support from body to lead.
4.1 Support Score Distributions
We find that significant fractions of lead and body claims are unsupported. Figure 3 plots (kernel density estimates of) lead and body claim support distributions for the PeopleProfiles dev split. We observe bimodality in both distributions, with high density around both full support (1.0) and no support (0.0). Indeed, 21.7% of lead claims are judged unsupported ($\leq 0$) by the body text and 29.5% of body claims by their cited source text(s). Despite the bimodality, a sizable proportion of claims do receive partial support (i.e. scores in (0,1)), including 32.6% of lead claims and 35.7% of body claims.
4.2 Unsupported Claims
It is natural to wonder what sorts of claims tend to receive support scores $\leq 0$. We find that the answer differs between lead and body claims. Table 2 presents several common categories of unsupported claims separately for the lead and body.
For lead claims, we observe that certain key facts—ubiquitous in lead sections—often do not receive any treatment in the body, including birth and death date and place (1st row), nicknames (2nd row), nationality (3rd row), and information about career status (4th row). In some cases (e.g. nicknames, nationality), the evidence itself may be difficult to obtain or arguably may not merit a citation. But in other cases (birth/death details, career status), the observed lack of body evidence seems to represent a departure from Wikipedia guidelines. [^12] E.g. “Birth and death places, if known, should be mentioned in the body of the article” (Wikipedia Manual of Style).
Categories of unsupported body claims are more varied, as article bodies themselves are more diverse. Here, some unsupported claims are ones that situate other facts that are more central to the article; others summarize (the import of) those facts (5th row). When the role of such claims is largely to contextualize or to recap ground that has been covered, the absence of supporting source material can perhaps be justified, though not always.
Hearsay or claims implying that some fact is disputed are another important case (6th row). In the example shown, although technically only one account of Washington’s opposition to flogging is required to establish that some accounts report that Washington opposed flogging, other accounts in the source for this claim say the opposite. Cases like this—where evidence is conflicting—can thus result in support labels of 0 or less.
Finally, claim decomposition (even after decontextualization) can occasionally result in what Jiang et al. (2024) term bleached claims—ones that are highly likely to be true, even independent of any evidence (7th row). For this reason, bleached claims are generally not ones that require evidence.
Figure 4: Distribution of overall evidence scores for PeopleProfiles dev split body evidence with mean- (blue) and product-based (orange) aggregation of body claim support scores for each evidence sentence.
4.3 Claim Support Propagation
@@
We annotate lead claim support given body evidence, but we can also consider how strong the support is for that evidence based on cited sources. We consider two methods of computing a support score for a body evidence sentence based on the support scores of the claims decomposed from it: taking the mean of those scores or taking their product. [^13] For product scores, we clip body claim scores <0<0 to 0. We can then compute an overall score for the set of evidence sentences for a given lead claim by applying the same aggregations (mean or product) across sentences.
First, we find that the majority of lead claims (82%) cannot be grounded in external source material via this method—i.e. by attempting to locate that material via its body evidence sentences. Importantly, this does not imply that evidence for all these lead claims is missing from those sources. Rather, the principal issue is that, for many lead claims, many of the body sentences in the evidence set for that claim do not bear a citation to a (publicly accessible) source. Absent a citation, we thus could not annotate claims in these body sentences for source support.
This finding underscores a key challenge for all NLP work focused on evidence attribution: real-world citation practices routinely deviate from the standard model (adopted here and elsewhere) in which each citation-worthy sentence is expected to bear supporting citations. In practice, a citation may be provided just once at the beginning of a paragraph (if it is clear that the whole paragraph concerns the same source) or lumped together with other citations only at the end of a paragraph. While such practices can obscure the mapping between claims and source material, and while they may diverge from documented standards for attribution (whether in Wikipedia or elsewhere), they are nonetheless commonplace and methods should be developed to better accommodate them.
Restricting ourselves to PeopleProfiles lead claims that can be grounded in sources, Figure 4 plots the distribution of overall evidence scores for lead claims from the dev split under both the mean (blue) and product (orange) aggregation strategies. We find very modest overall scores in both cases (an average of 0.41 for mean and 0.35 for product).
5 Evidence Retrieval
We now turn our focus from the extent of evidential support for a target claim to retrieval of find-grained evidence for such a claim. We consider two claim-centric evidence retrieval tasks: